# Transparent Graphene-Superconductor Interfaces: Quantum Hall and Zero Field Regimes

### v1

@@ I.1 Motivation and Context
The interface between graphene and a superconductor has attracted great interest for nearly two decades from several points of view (for reviews see Refs. [1, 2]). First, there has been a long-standing desire for semiconductor-superconductor structures that are gateable and so controllable [3, 4], leading potentially to a variety of interesting quantum devices [5, 6, 7, 8]. By encapsulating a graphene sheet between two cladding layers (typically BN), such a high mobility and easily gateable structure can be made from graphene [9, 10, 11, 12, 13].
@@
Theoretical interest in the QH/superconductor system predates the recent experiments [40, 41, 42, 43, 44, 45, 46, 47, 48, 49] but has certainly increased because of them. In particular, recent papers [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63] addressed, for instance, interference among interface modes, effects of sample geometry and disorder, and the roles of dissipation and supercurrent in QH/superconductor hybridization, though not necessarily in the context of graphene.
Our goal in this work is to understand the properties of infinite, uniform graphene/superconductor interfaces in the QH regime, as well as at zero magnetic field. Because the experimental interfaces are thought to be highly transparent [29, 26], we consider interfaces in which the electrons in graphene are coupled strongly to those in the superconductor: there should be no interfacial barrier and the wavefunctions should match well (explained further below). We refer to this situation loosely by saying the interface is “transparent”. For such interfaces, we find that the special properties of graphene lead to new behavior.
@@ I.2 Summary of Main Results
We present results for a clean interface between a zigzag edge of graphene in the QH regime and a superconductor, using both continuum and tight-binding models. In both cases we solve the single-particle Bogoliubov-de Gennes equations (BdG) that describe inhomogeneous, mean-field superconductors [64, 1]. In the continuum model, we match wavefunctions between a Dirac spectrum material and one with parabolic dispersion with a large Fermi energy. In the tight-binding approach, we stitch together a honeycomb lattice for the graphene with a square lattice near half filling for the superconductor. The two main assumptions are that (i) there is no disorder in either material or at the interface and (ii) the fields in the BdG equations (the pair potential, one-body potential, and magnetic field) all change abruptly at the graphene-superconductor interface.
@@
Figure 1: Transparent interface between a superconductor ($y>0$) and quantum Hall graphene in the LLL ($y<0$) in the continuum model. (a) Wavefunction of the interface mode, $|\psi(y)|^{2}$, in the $K$ valley: the electron and hole components are plotted as positive and negative numbers, respectively, and in graphene the sublattice A (B) component is green (blue). In graphene the electron and hole components have the same magnitude, indicating perfect Andreev reflection, but the sublattices are exchanged. (b) Electron-hole hybridization expressed as the fractional electron content in graphene [Eq. (1)] as a function of the LLL edge state filling. Shown for three values of the superconductor Fermi energy [cf. Eq. (17) for the definition of $\tilde{v}^{2}$]. Excellent hybridization is obtained throughout the LLL (except very near the Dirac point). (c) Dispersion of the chiral edge mode ($E<\Delta$) for different strengths of the interface barrier (varying $Z=V_{0}/\hbar v$ to color-coded values). The spectrum evolves from the terminated-lattice edge mode on the right to the fully-hybridized interface mode centered on the $K$ point ($kl_{B}=0$). (The discontinuities in the spectrum are unrelated to the interface state.) For a transparent interface, then, the spectrum is valley degenerate. (Zigzag edge using standard parameters, see Sec. II.4.)
@@
First, for a transparent interface, we find strong Andreev reflection—in the quantum Hall regime as well as at zero magnetic field. Current continuity at the interface implies that the ratio of velocities, $v/v_{FS}$, is an important parameter (where $v$ is the velocity at the Dirac point in graphene and $v_{FS}$ is the Fermi velocity in the superconductor) and that one should have $v/v_{FS}\!\sim\!1$ for good matching. Since $v\!\approx\!10^{6}$ m/sec independent of the graphene chemical potential, achieving this condition is straightforward—unlike for a semiconductor with parabolic dispersion in which the velocity is typically small. We find a wide range of parameters that yield excellent Andreev reflection.
In the QH regime, strong Andreev reflection is illustrated in Fig. 1(a) which shows the transverse wavefunction of the quasi-one-dimensional state at the graphene-superconductor interface in the lowest Landau level (LLL). On the graphene side ($y<0$), apart from a swap of sublattices, the electron and hole densities are nearly identical, showing that Andreev reflection is strong. Further discussion of the transverse wavefunction appears below in Sec. III.2, including for non-ideal cases in V.
@@
The paper is organized as follows. In Sec. II, we introduce the specific continuum and tight-binding models, discuss the approximations, and give our typical values for the model parameters. Sec. III focuses on strong Andreev reflection for transparent interfaces, first at $B=0$ and then in the QH regime. In Sec. IV, we discuss the dispersion of the interface modes and the question of valley degeneracy for both perfect and imperfect interfaces. We argue that valley degeneracy, strong Andreev reflection, and strong coupling are naturally connected (Sec. IV.4). The dependence on properties of the superconductor, in particular its chemical potential, are discussed in Sec. V. Though the spin degree of freedom is unimportant in most of our work, in Sec. VI we discuss the Zeeman effect, showing how it breaks perfect hybridization but does not break valley degeneracy. Implications for Majorana modes are discussed. Finally, our conclusions are in Sec. VII, followed by several appendices.
@@ II.2 Tight-Binding Model of the Interface
Since a zigzag edge is thought to be a good model for a generic graphene edge [73, 74, 53], we consider a graphene half-sheet with a zigzag edge (Fig. 2). The graphene half-sheet is translationally invariant in the $x$ direction with period $a$, the lattice constant of graphene (next-nearest-neighbor distance). Our convention is that the sites on the terminal row of the zigzag edge are at $y=0$ and belong to the A sublattice. For the magnetic field, we work in the Landau gauge with $\mathbf{A}(x,y)=By\hat{x}\,\theta(-y)$ where $\theta(y)$ is the step function [thus $\mathbf{B}=-B\hat{z}\,\theta(-y)$]. With these choices, $\Delta$ is constant and real.
@@ II.3 Continuum Model of the Interface
The relative magnitude of graphene’s Dirac point velocity to the perpendicular velocity in the superconductor is clearly important for current continuity, Eq. (15). Their ratio is a key parameter below, and so we define
@@ II.4 Regime Studied and Parameter Values
In the tight-binding model, the hopping matrix element in both the graphene and superconductor is $t$. Our standard parameter values are $t_{NS}\!=\!t$, $E_{L}/t\!=\!0.12$, $\mu_{\textrm{gr}}/E_{L}\!=\!1/4$, and $\Delta/E_{L}\!=\!1/10$ [80]. The magnetic length and magnetic field corresponding to this $E_{L}$ are $l_{B}/a\!=\!10.2$ and $Ba^{2}/\Phi_{0}\!=\!0.0015$ (or a flux of $0.0013\,\Phi_{0}$ per unit cell, where $a$ is the lattice constant of graphene). In the superconductor, $\mu_{S}$ is typically around half filling, yielding $\mu_{S}/t\!\sim\!4$. (The “optimal values” of $\mu_{S}/t$ that we use are $4.5$, $5.0$, and $0.2$ for the dense, sparse, and honeycomb lattices, respectively.) Thus the Fermi energy in graphene is much smaller than that in the superconductor: $\mu_{\textrm{gr}}/\mu_{S}\sim 0.01$.
@@ III.1 $B=0$: Robust Andreev Reflection and the Dirac Point Velocity
This is very different from Andreev reflection in a semiconductor with parabolic dispersion. In that case, the big mismatch between the velocities in the semiconductor and the superconductor suppresses Andreev reflection strongly. Indeed, with the same simplifications as for graphene above, the result for a clean semiconductor-superconductor interface is
@@
where $N$ denotes the semiconductor [68, 69, 70, 71]. Thus, the conical valleys in graphene allow much better connection to superconductors than for semiconductors with comparable density; indeed, in graphene there is the possibility of near perfect Andreev reflection.
@@ III.2.1 QH Edge States on a Terminated Lattice
When $\overline{y}$ approaches the edge, boundary conditions come into play. In our geometry the boundary condition is that $\psi$ on the B sublattice must be zero (i.e. the absent next row in the lattice). Since the $K^{\prime}$ valley states are non-zero only on A sites, they are not modified: their energy remains strictly zero.
@@ III.2.2 Interface wave function: Continuum model
Three length scales are visible in the wavefunction. (i) The magnetic length, $l_{B}$, sets the scale of variation in the quantum Hall state. (ii) The Fermi wavelength in the superconductor sets the scale for the rapid oscillations at $y>0$. (iii) The scale of the slowly decaying envelope in the superconductor is given by the coherence length $\propto 1/\Delta$  [^2] More precisely, for quantities (ii) and (iii) one must take the projection perpendicular to the interface [see Eq. (10)]: for (ii), the period of the oscillations is $2\pi/q^{\prime}\!\approx\!2\pi/q_{y}$, and for (iii), the coherence length perpendicular to the interface is $\xi_{y}\equiv 1/q^{\prime\prime}\!\approx\!\hbar v_{y}/\Delta$ which, for small $\Delta$, is much larger than the period. The rapid oscillations of the electron and hole wavefunctions in the superconductor are exactly out of phase, as expected for a mid-gap state ($E=0$ implies $\gamma=-i$)..
The filling of the edge state, $\mu_{\textrm{gr}}/E_{L}$, plays little role in the matching. Indeed, for a suitable superconductor for which $\widetilde{v}^{2}\approx 1$, the degree of electron-hole hybridization is essentially independent of $\mu_{\textrm{gr}}$. The fractional electron content in graphene, $f_{e}$ in Eq. (1), is one way to quantify the $e$-$h$ hybridization. Fig. 1(b) shows $f_{e}$ in the LLL as $\mu_{\textrm{gr}}$ varies from the Dirac point to the threshold for the first Landau level, $E_{L}$. For $\widetilde{v}^{2}=1$, hybridization is perfect, $f_{e}=1/2$, throughout the LLL.
@@ IV.1 QH Terminated-Lattice Spectrum
The spectrum with both electrons and holes is shown in Fig. 6 [89]. In the continuum solution, panels (a)-(b), both the electron mode at the chemical potential in the $K$ valley and the corresponding hole mode in $K^{\prime}$ are clearly seen. Since the LLL $K^{\prime}$ valley states are not modified by the termination of the lattice (they are non-zero only on A sites), their energy remains zero: they appear as horizontal blue lines in panel (b). Due to this flat band, while there is no hole state in the $K$ valley at the chemical potential (no red line crosses $E=0$), the electron and hole solutions do intersect at $E=\mu_{\textrm{gr}}$.
For the tight-binding results, panel (c), periodicity in the $x$ direction leads to a Brillouin zone of size $2\pi/a$. We show the interval $k\in[0,2\pi/a]$ in which the Dirac points are $Ka\!=\!2\pi/3$ and $K^{\prime}a\!=\!4\pi/3$. To focus on the $y\!=\!0$ edge of the nanoribbon, the width of the nanoribbon, $W$, is chosen so that the spectrum of the states on the other edge is clearly separated: $W\lesssim l_{B}^{2}(4\pi/3a)$  [^4] The choice $W\!\lesssim l_{B}^{2}(4\pi/3a)$ avoids having the spectrum of the bulk states wrap around the Brillouin zone, which would then overlap the spectrum of the edge states near $y\!=\!0$..
@@ IV.2 Implications for Andreev Reflection
At first sight, the case for a strong proximity effect appears hopeless: we should look for electron and hole states that (i) have the same crystal momentum (translational invariance), (ii) overlap spatially in the transverse direction (local interaction), and (iii) differ in energy by $\lesssim\Delta$ (matrix element of superconductivity that mixes electrons and holes). These criteria are never satisfied in Fig. 6: there is no crossing of the electron and hole modes at the chemical potential. Such a crossing is achieved only through fine tuning to $\mu_{\textrm{gr}}/E_{L}=1$.
@@
However, we find that there is a second scenario that leads to a strong proximity effect in the clean limit, namely if the interface is transparent. Fig. 7 shows the evolution of the spectrum as the graphene-superconductor hopping matrix element, $t_{NS}$, increases. (In order to clearly see the LLL states at $E=\pm\mu_{\textrm{gr}}$, a large value of $\Delta$ is used, $\Delta/\mu_{\textrm{gr}}\!=\!2$, so that they are not hidden by the superconductor’s continuum.) For small $t_{NS}$ (green line), an avoided crossing appears in the $K$ valley at $E=\mu_{\textrm{gr}}$ where the electron and hole states of the terminated-lattice spectrum cross. The avoided crossing indicates electron-hole hybridization. As $t_{NS}$ grows, the spectrum changes throughout the Brillouin zone and goes well beyond a simple avoided crossing: one state is pushed to higher energy while the other becomes a CAEM with antisymmetry about the $K$ point (black). The large hopping matrix element at a transparent interface can couple many Landau levels, $t_{NS}\!\gg\!E_{L}$ (in Fig. 7, $t/E_{L}\!=\!8$), and completely reorganizes the electron and hole states. In this reorganization, the small matrix element of superconductivity readily causes electron-hole hybridization.
We conclude that while a superconductor cannot induce a proximity effect on a clean zigzag QH edge through tunneling—the dead end applies when $t_{NS}\!\ll\!t$—a strong proximity effect can readily occur when $t_{NS}\!\sim\!t$ so that the interface is transparent. The induced pairing correlations are quite different in these two scenarios. In the first (disorder, tunneling), sublattice-valley locking implies that the graphene sublattice structure of the electrons and holes is the same. In the second scenario (clean, transparent), the electrons and holes are on opposite sublattices, as in Figs. 1(a) and 5.

### v2

@@ I.1 Motivation and Context
The interface between graphene and a superconductor has attracted great interest for nearly two decades from several points of view (for reviews see Refs. [1, 2]). First, there has been a long-standing desire for semiconductor-superconductor structures that are gateable and so controllable [3, 4], leading potentially to interesting quantum devices [5, 6, 7, 8]. By encapsulating graphene between two cladding layers (typically BN), such a high mobility and gateable structure can be made from graphene [9, 10, 11, 12, 13].
@@
Theoretical interest in the QH/superconductor system predates the recent experiments [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50] but has certainly increased because of them. In particular, recent papers [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64] addressed, for instance, interference among interface modes, effects of sample geometry and disorder, and the roles of dissipation and supercurrent in QH/superconductor hybridization, though not necessarily in the context of graphene.
Our goal in this work is to understand the properties of infinite, uniform graphene/superconductor interfaces in the QH regime. Because the experimental interfaces are thought to be highly transparent [29, 26], we consider interfaces in which the electrons in graphene are coupled strongly to those in the superconductor: there should be no interfacial barrier and the wavefunctions should match well (explained further below). We refer to this situation loosely by saying the interface is “transparent”. For such interfaces, we find that the special properties of graphene lead to new behavior.
@@ I.2 Summary of Main Results
We present results for a clean interface between a zigzag edge of graphene in the QH regime and a superconductor, using both continuum and tight-binding models. In both cases we solve the single-particle Bogoliubov-de Gennes equation (BdG) that describes inhomogeneous, mean-field superconductors [65, 1]. In the continuum model, we match wavefunctions between a Dirac spectrum material and one with parabolic dispersion with a large Fermi energy. In the tight-binding approach, we stitch together a honeycomb lattice for the graphene with a square lattice near half filling for the superconductor. The two main assumptions are that (i) there is no disorder in either material or at the interface and (ii) the fields in the BdG equation (the pair potential, one-body potential, and magnetic field) all change abruptly at the graphene-superconductor interface.
@@
Figure 1: Transparent interface between a superconductor ($y>0$) and quantum Hall graphene in the LLL ($y<0$) in the continuum model. (a) Wavefunction of the interface mode, $|\psi(y)|^{2}$, in the $K$ valley: the electron and hole components are plotted as positive and negative numbers, respectively, and in graphene the sublattice A (B) component is green (blue). In graphene the electron and hole components have the same magnitude, indicating perfect Andreev reflection, but the sublattices are exchanged. (b) Electron-hole hybridization expressed as the fractional electron content in graphene [Eq. (1)] as a function of the LLL edge state filling ($E_{L}\equiv\sqrt{2}\hbar v/l_{B}$). Shown for three values of the superconductor Fermi energy [cf. Eq. (17) for the definition of $\tilde{v}^{2}$]. Excellent hybridization is obtained throughout the LLL (except very near the Dirac point). (c) Dispersion of the chiral edge mode ($E<\Delta$) for different strengths of the interface barrier (varying $Z=V_{0}/\hbar v$ to color-coded values). The spectrum evolves from the terminated-lattice edge mode on the right to the fully-hybridized interface mode centered on the $K$ point ($kl_{B}=0$). (The discontinuities in the spectrum are unrelated to the interface state [70].) For a transparent interface, then, the spectrum is valley degenerate. (Zigzag edge, $\Delta=E_{L}/10$, for other parameters see Sec. II.4.)
@@
First, for a transparent interface, we find strong Andreev reflection—in the quantum Hall regime as well as at zero magnetic field. Current continuity at the interface implies that the ratio of velocities, $v/v_{FS}$, is an important parameter (where $v$ is the velocity at the Dirac point in graphene and $v_{FS}$ is the Fermi velocity in the superconductor) and that one should have $v/v_{FS}\!\sim\!1$ for good matching. Since $v\!\approx\!10^{6}$ m/sec independent of the graphene chemical potential, achieving this condition is straightforward—unlike for a semiconductor with parabolic dispersion in which the velocity is typically small [71]. We find a wide range of parameters that yield excellent Andreev reflection.
In the QH regime, strong Andreev reflection is illustrated in Fig. 1(a) which shows the transverse wavefunction of the quasi-one-dimensional state at the graphene-superconductor interface in the lowest Landau level (LLL). On the graphene side ($y<0$), apart from a swap of sublattices, the electron and hole densities are nearly identical, showing that Andreev reflection is strong. Further discussion of the transverse wavefunction appears in Sec. III.2, including for non-ideal cases in V.
@@
The paper is organized as follows. In Sec. II, we introduce the specific continuum and tight-binding models, discuss the approximations, and give our typical values for the model parameters. Sec. III focuses on strong Andreev reflection for transparent interfaces, first at $B=0$ and then in the QH regime. In Sec. IV, we discuss the dispersion of the interface modes and the question of valley degeneracy for both perfect and imperfect interfaces. We argue that valley degeneracy, strong Andreev reflection, and strong coupling are naturally connected (Sec. IV.4). The dependence on properties of the superconductor, in particular its chemical potential, are discussed in Sec. V. Though the spin degree of freedom is unimportant in most of our work, in Sec. VI we discuss the Zeeman effect, showing how it breaks perfect hybridization but does not break valley degeneracy. Implications for Majorana modes are discussed. Finally, our conclusions are in Sec. VII: they include a discussion connecting these infinite-interface results (in particular, $f_{e}$) to experiments, which are typically performed in nanostructures.
@@ II.2 Tight-Binding Model of the Interface
Since a zigzag edge is thought to be a good model for a generic graphene edge [75, 76, 54], we consider a graphene half-sheet with a zigzag edge (Fig. 2) [77]. The graphene half-sheet is translationally invariant in the $x$ direction with period $a$, the lattice constant of graphene (next-nearest-neighbor distance). Our convention is that the sites on the terminal row of the zigzag edge are at $y=0$ and belong to the A sublattice. For the magnetic field, we work in the Landau gauge with $\mathbf{A}(x,y)=By\hat{x}\,\theta(-y)$ where $\theta(y)$ is the step function [thus $\mathbf{B}=-B\hat{z}\,\theta(-y)$]. With these choices, $\Delta$ is constant and real.
@@ II.3 Continuum Model of the Interface
We emphasize that current continuity is a condition on the velocity at the interface in the two materials. The relative magnitude of graphene’s Dirac point velocity to the perpendicular velocity in the superconductor is clearly important in Eq. (15). Their ratio is a key parameter below, and so we define
@@ II.4 Regime Studied and Parameter Values
In the tight-binding model, the hopping matrix element in both the graphene and superconductor is $t$. Our standard parameter values are $t_{NS}\!=\!t$, $E_{L}/t\!=\!0.12$, $\mu_{\textrm{gr}}/E_{L}\!=\!1/4$, and $\Delta/E_{L}\!=\!1/10$ [82]. The magnetic length and magnetic field corresponding to this $E_{L}$ are $l_{B}/a\!=\!10.2$ and $Ba^{2}/\Phi_{0}\!=\!0.0015$ (or a flux of $0.0013\,\Phi_{0}$ per unit cell), where $a$ is the lattice constant of graphene. In the superconductor, $\mu_{S}$ is typically around half filling, yielding $\mu_{S}/t\!\sim\!4$. (The “optimal values” of $\mu_{S}/t$ that we use are $4.5$, $5.0$, and $0.2$ for the dense, sparse, and honeycomb lattices, respectively.) Thus the Fermi energy in graphene is much smaller than that in the superconductor: $\mu_{\textrm{gr}}/\mu_{S}\sim 0.01$.
@@ III.1 $B=0$: Robust Andreev Reflection and the Dirac Point Velocity
This is very different from Andreev reflection in a semiconductor with parabolic dispersion. In that case, the big mismatch between the velocities in the semiconductor and the superconductor [71] suppresses Andreev reflection strongly. With the same simplifications as for graphene above, the result for a clean semiconductor-superconductor interface is
@@
where $N$ denotes the semiconductor [69, 72, 73, 74]. Thus, the conical valleys in graphene allow much better connection to superconductors than for semiconductors with comparable density: the honeycomb lattice provides a natural decoupling of the Fermi velocity (appearing in current continuity) from the Fermi wavevector. As a result, in graphene there is the possibility of near perfect Andreev reflection.
@@ III.2.1 QH Edge States on a Terminated Lattice
When $\overline{y}$ approaches the $y\!=\!0$ edge, boundary conditions come into play. In our geometry the boundary condition is that $\psi$ on the B sublattice must be zero (i.e. the absent next row in the lattice). Since the $K^{\prime}$ valley states are non-zero only on A sites, they are not modified: their energy remains strictly zero.
@@ III.2.2 Interface wave function: Continuum model
Three length scales are visible in the wavefunction. (i) The magnetic length, $l_{B}$, sets the scale of variation in the quantum Hall state. (ii) The Fermi wavelength in the superconductor sets the scale for the rapid oscillations at $y>0$. (iii) The scale of the decaying envelope in the superconductor is the coherence length $\propto 1/\Delta$  [^2] More precisely, for quantities (ii) and (iii) one must take the projection perpendicular to the interface [see Eq. (10)]: for (ii), the period of the oscillations is $2\pi/q^{\prime}\!\approx\!2\pi/q_{y}$, and for (iii), the coherence length perpendicular to the interface is $\xi_{y}\equiv 1/q^{\prime\prime}\!\approx\!\hbar v_{y}/\Delta$ which, for small $\Delta$, is much larger than the period. The rapid oscillations of the electron and hole wavefunctions in the superconductor are exactly out of phase, as expected for a mid-gap state [$E=0$ implies $\gamma=-i$ in Eq. (35)]..
The filling of the edge state, $\mu_{\textrm{gr}}/E_{L}$, plays little role in the matching. Indeed, for a suitable superconductor for which $\widetilde{v}^{2}\approx 1$, the degree of electron-hole hybridization is essentially independent of $\mu_{\textrm{gr}}$. The fractional electron content in graphene, $f_{e}$ in Eq. (1), is one way to quantify the $e$-$h$ hybridization. Fig. 1(b) shows $f_{e}$ in the LLL as $\mu_{\textrm{gr}}$ varies from the Dirac point to the threshold for the first Landau level, $E_{L}$. For $\widetilde{v}^{2}=1$, hybridization is perfect, $f_{e}=1/2$, throughout the LLL. Thus, as in the $B\!=\!0$ case, we see that the decoupling of the Fermi velocity and conduction electron density enables strong Andreev reflection.
@@ IV.1 QH Terminated-Lattice Spectrum
The spectrum with both electrons and holes is shown in Fig. 6 [91]. In the continuum solution, panels (a)-(b), both the electron mode at the chemical potential in the $K$ valley and the corresponding hole mode in $K^{\prime}$ are clearly seen. Since the LLL $K^{\prime}$ valley electron states are not modified by the termination of the lattice (they are non-zero only on A sites), their energy remains zero: they appear as a horizontal blue line in panel (b). Due to this flat band, while there is no hole state in the $K$ valley at the chemical potential (no red line crosses $E=0$), the electron and hole solutions do intersect at $E=\mu_{\textrm{gr}}$.
For the tight-binding results, panel (c), periodicity in the $x$ direction leads to a Brillouin zone of size $2\pi/a$. We show the interval $k\in[0,2\pi/a]$ in which the Dirac points are $Ka\!=\!2\pi/3$ and $K^{\prime}a\!=\!4\pi/3$. To focus on the $y\!=\!0$ edge of the nanoribbon, the width of the nanoribbon, $W$, is chosen so that the spectrum of the states on the other edge is clearly separated: $W\lesssim l_{B}^{2}(4\pi/3a)$  [^4] The choice $W\!\lesssim l_{B}^{2}(4\pi/3a)$ avoids having the spectrum of the bulk states wrap around the Brillouin zone, which would then overlap the spectrum of the edge states near $y\!=\!0$.. With the gauge that we use, the QH states with guiding center coordinate at the interface, $\overline{y}\!=\!0$, correspond to $k=K\!=\!2\pi/3a$ and $k=K^{\prime}\!=\!4\pi/3a$.
@@ IV.2 Implications for Andreev Reflection
At first sight, the case for a strong proximity effect appears hopeless: we should look for electron and hole states that (i) have the same crystal momentum (translational invariance), (ii) overlap spatially in the transverse direction (local interaction), and (iii) differ in energy by $\lesssim\Delta$ (matrix element of superconductivity that mixes electrons and holes). These criteria are never satisfied in Fig. 6: there is no crossing of the electron and hole modes at the chemical potential.
@@
However, we find that there is a second scenario that leads to a strong proximity effect in the clean limit, namely if the interface is transparent. Fig. 7 shows the evolution of the spectrum as the graphene-superconductor hopping matrix element, $t_{NS}$, increases. (In order to clearly see the LLL states at $E=\pm\mu_{\textrm{gr}}$, a large value of $\Delta$ is used, $\Delta/\mu_{\textrm{gr}}\!=\!2$, so that they are not hidden by the superconductor’s continuum.)
In the disconnected case (blue dots), there is a crossing in each valley between three states: in the $K$ valley, for instance [see also Fig. 6(a)], the rapidly dispersing terminated-lattice edge state intersects two non-dispersive hole states at $E\!=\!\mu_{\textrm{gr}}$ (horizontal blue line). These holes correspond to the degenerate electron states at $E\!=\!-\mu_{\textrm{gr}}$ in the $K^{\prime}$ valley, whose nature is discussed in the graphene nanoribbon literature [84, 85, 86, 76, 93]. We learn that one of the non-dispersive hole states is the normal $K$-valley LLL hole state, which is not dispersive at the edge because it has weight on only the A sublattice while the boundary condition is on the B sublattice. The other non-dispersive hole state is necessarily far from the $y\!=\!0$ interface: depending on the nanoribbon width, it is either the $K^{\prime}$ valley LLL state with guiding center deep in the bulk or the anomalous zigzag-nanoribbon state localized on the far edge of the nanoribbon.
For small $t_{NS}$ (green), an avoided crossing appears between the two states near the interface: the electron edge state and the normal $K$-valley hole state avoid crossing, while the state far from $y\!=\!0$ remains unchanged. The avoided crossing indicates electron-hole hybridization.
As $t_{NS}$ grows, the spectrum changes throughout the Brillouin zone and goes well beyond a simple avoided crossing: one state is pushed to higher energy while the other becomes a CAEM with antisymmetry about the $K$ point (black). The large hopping matrix element at a transparent interface can couple many Landau levels, $t_{NS}\!\gg\!E_{L}$ (in Fig. 7, $t/E_{L}\!=\!8$), and reorganizes the electron and hole states, leading to electrons and holes at nearly the same energy. The small pairing matrix element $\Delta$ then readily causes electron-hole hybridization. The big reorganization of the electron and hole states implies that their original non-dispersive character due to special zigzag-nanotube characteristics is not important [94]: the CAEM are produced by a combination of large $t_{NS}$ and small $\Delta$ and cannot be thought of as hybridization between only two terminated-lattice states.
We conclude that while a superconductor cannot induce a proximity effect on a clean zigzag QH edge through tunneling—the dead end applies when $t_{NS}\!\ll\!t$—a strong proximity effect can readily occur when $t_{NS}\!\sim\!t$ so that the interface is transparent. Notably, the induced pairing correlations are quite different in these two scenarios. In the first (disorder, tunneling), sublattice-valley locking implies that the graphene sublattice structure of the electrons and holes is the same. In the second scenario (clean, transparent), the electrons and holes are on opposite sublattices, as in Figs. 1(a) and 5.